OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 1) greater than or equal to -12? x greater than or equal to -5 x less than or greater to -5 -5 less than or greater to x A number line with a closed circle on negative 5 and shading to the right

OpenStudy (anonymous):

\[2x-2\ge x\] =\[2x \ge x+2\] =\[2x-x \ge2\]= \[x \ge2\]

OpenStudy (ashleyisakitty):

Next time post in mathematics.